OpenStudy (anonymous):

write 0.5 (repeating) as a fraction simplify if needed

OpenStudy (anonymous):

\[\frac{5}{9}\]

OpenStudy (anonymous):

\[\bar{.1}=\frac{1}{9}\] \[\bar{.2}=\frac{2}{9}\] \[\bar{.3}=\frac{3}{9}\] \[\bar{.4}=\frac{4}{9}\]

OpenStudy (anonymous):

thanks